Three types of concert tickets are being sold. The most expensive tickets are twice expensive as the balcony tickets. Balcony tickets are $10 less than the main floor tickets and $30 less than the most expensive. Determine the price of each ticket.
Answer by stanbon(75887)

Three types of concert tickets are being sold. The most expensive tickets are twice expensive as the balcony tickets. Balcony tickets are $10 less than the main floor tickets and $30 less than the most expensive. Determine the price of each ticket.
Equations:
e = 2b
b = m-10
b = e-30
---------
Substitute for "e" and solve for "b"::
b = 2b-30
b = 30 (price of balconey tickets)
----
e = 2b = 60 (price of most expensive tickets)
----
m = b+10 = 40 (price of main floor tickets)
================
Cheers,
Stan H.
================